# Preparations for Quantum Computing in Hadron Physics

### v1

@@ 5.1.3 $n$-particle operators
Generalizing section 5.1.2, $n$-register operators are constructed by means of projectors (control), $\mathfrak{C}_{ij}=|i\rangle\langle j|$, identities $\mathfrak{i}$ and set/scrap operators $\mathfrak{s}^{\dagger}_{p}/\mathfrak{s}_{p}$. The operator that activates the first register is
@@
$a^{(n)\dagger}_{p,1}{\left|{\Omega}\right\rangle}$ $=$ $\left({\text{$\mathfrak{C}_{{0}{0}}$}}{\left|{0}\right\rangle}_{P/A}{\text{$\mathfrak{i}$}}{\left|{0...0}\right\rangle}\right)_{n}...\left({\text{$\mathfrak{C}_{{1}{0}}$}}{\left|{0}\right\rangle}_{P/A}\mathfrak{s}^{\dagger}_{p}{\left|{0...0}\right\rangle}\right)_{1}$ (41)
$=$ $\left({\left|{0}\right\rangle}_{P/A}{\left|{0...0}\right\rangle}\right)_{n}...\left({\left|{1}\right\rangle}_{P/A}{\left|{p}\right\rangle}\right)_{1}\ .$
The most efficient way to implement the symmetrization for more than two registers is to proceed in a stepwise fashion, assuming previously symmetrized n1n-1 registers and requiring the newly filled register to also be in a symmetric wavefunction with the remaining ones. For instance, the three particle symmetrizer needed upon adding a third particle is
$\mathcal{S}_{3}\rightarrow\mathcal{S}_{3\leftarrow 2}\equiv\frac{1}{\sqrt{3}}\left({\text{$\mathbb{I}$}}^{\otimes 3}+\mathcal{P}_{32}\otimes{\text{$\mathbb{I}$}}_{1}+\mathcal{P}_{31}\otimes{\text{$\mathbb{I}$}}_{2}\right)\ ,$ (42)
and, the $n$-particle symmetrizer [^2] This symmetrizer is not idempotent, but satisfies $S^{2}_{n\leftarrow n-1}=\sqrt{n}\ S_{n\leftarrow n-1}$ instead.
@@
$a^{(n)\dagger}_{q}=\sum^{n}_{i=1}a^{(n)\dagger}_{q,i}=\sum^{n}_{i=1}\mathcal{S}_{i\leftarrow(i-1)}\cdot{\text{$\mathbb{P}^{({n-i})}_{{0}}$}}\otimes\left({\text{$\mathfrak{C}_{{1}{0}}$}}\otimes\mathfrak{s}^{\dagger}_{q}\right)_{i}\otimes\,{\text{$\mathbb{P}^{({i-1})}_{{i-1}}$}}$, (44)
@@
which would be impossible to satisfy with operators acting on single registers (non-trivially) if the memory was not explicitly symmetrized.
@@
. With this definitions we can any operator written in terms of creators and annihilators, for the number operator we have
@@
$=\sum_{j}\,j\,{\text{$\mathbb{P}^{({n})}_{{j}}$}}$., (48)
where in the last line we assumed a symmetric memory so that $\mathcal{S}_{j\leftarrow j-1}$ (c.f. 43) simplify to $\sqrt{j}$.
@@ 5.2 Fermions
Because the conceptual architecture is very similar, we only briefly touch on anticommuting operators. We change the recursive step-symmetrizer operators to step-antisymmetrizers
${\text{$\mathcal{A}_{{n}\leftarrow{n}-1}$}}=\frac{1}{\sqrt{n}}\left({\text{$\mathbb{I}$}}^{\otimes n}-\mathcal{P}_{n(n-1)}-\mathcal{P}_{n(n-2)}-...-\mathcal{P}_{n2}-\mathcal{P}_{n1}\right)$ (49)
@@
$b^{(n)\dagger}_{q}=\sum^{n}_{j=1}b^{(n)\dagger}_{q,j}=\sum^{n}_{j=1}{\text{$\mathcal{A}_{{j}\leftarrow{j}-1}$}}\cdot{\text{$\mathbb{P}^{({n-j})}_{{0}}$}}\otimes\underbrace{\left({\text{$\mathfrak{C}_{{1}{0}}$}}\otimes\mathfrak{s}^{\dagger}_{q}\right)_{j}}\otimes\,{\text{$\mathbb{P}^{({j-1})}_{{j-1}}$}}$ (50)
@@
$b^{(n)}_{q}=\sum^{n}_{j=1}b^{(n)}_{q,j}=\sum^{n}_{j=1}{\text{$\mathbb{P}^{({n-j})}_{{0}}$}}\otimes\left({\text{$\mathfrak{C}_{{0}{1}}$}}\otimes{\text{$\mathfrak{s}_{{p}}$}}\right)_{j}\otimes\,{\text{$\mathbb{P}^{({j-1})}_{{j-1}}$}}\cdot\mathcal{A}_{j\leftarrow j-1}$ (51)
They now fulfill the required anticommutation relations
@@ 5.3 Quantum evolution
For interaction terms that change the values of momenta we need to define a discretization of $p$; for instance, the equidistant grid
$\left\{p_{min},...,p_{-1},p_{0}=0,p_{1},...,p_{max}\right\}$ (55)
@@
Taking into account the structure of the operators in Eq. (50), and the fact that for particle-conserving interactions the Fock operators are diagonal, and therefore $e^{A+B}=e^{A}e^{B}$ the evolution operator can be expressed as:
$\mathcal{U}^{f}_{22}(\Delta t)={\text{$\mathbb{P}^{({n})}_{{0}}$}}+{\text{$\mathbb{P}^{({n})}_{{1}}$}}+\sum^{n-1}_{j=2}{\text{$\mathbb{P}^{({n-1-j})}_{{0}}$}}\otimes\left\{\left({\text{$\mathfrak{C}_{{1}{1}}$}}\right)^{\otimes j}\right\}\left\{\prod^{j}_{k=1}\prod^{j+1}_{l=1+k}\mathfrak{U}_{22,(l,k)}(\Delta t)\right\},$ (57)
with
$\frac{i}{2\Delta t}\log\left(\mathfrak{U}_{22,l.k}(\Delta t)\right)=$
$\sum_{\xi}\lambda_{\xi}\left\{\!\!\left(\sum_{p}{\text{$\mathfrak{s}^{\dagger}_{{p+\xi\Delta}}$}}{\text{$\mathfrak{s}_{{p}}$}}\right)_{l}\!\!\otimes\!\!\left(\sum_{q}{\text{$\mathfrak{s}^{\dagger}_{{q}}$}}{\text{$\mathfrak{s}_{{q+\xi\Delta}}$}}\right)_{k}\!\!+\!\!\left(\sum_{q}{\text{$\mathfrak{s}^{\dagger}_{{q}}$}}{\text{$\mathfrak{s}_{{q+\xi\Delta}}$}}\right)_{l}\!\!\otimes\!\!\left(\sum_{p}{\text{$\mathfrak{s}^{\dagger}_{{p+\xi\Delta}}$}}{\text{$\mathfrak{s}_{{p}}$}}\right)_{k}\!\!\right\}\ \ \$(58)
which is an auxiliary “momentum exchanger” among registers (in which all identity operators have been omitted to alleviate notation).
When considering Hamiltonian interaction terms that alter the number of particles, we require a dynamical memory, which changes the number of active registers.
Similarly, we can define tadpoles, emission or absorption of particles, etc. For instance, the following operator stands for the evolution associated to the Hamiltonian term corresponding to the quark-gluon vertex
@@ 6 Gauge fixing
The gauge redundancy in Yang-Mills theory or Chromodynamics forces one to select a gauge when fixing the physical states in any Hamiltonian formulation. First and foremost we discuss equal–time quantization in subsection 6.1. Then we turn to light–front quantization in subsection 6.2.
With a bird’s eye view, unless one fixes Coulomb gauge which is a complete gauge fixing with an entirely physical Fock space but a convoluted Hamiltonian necessitating approximation [117, 118, 119, 120], the computer will often be approximating a redundant Fock space on which Gauss’s law needs to be employed to project over a physical instance. The later evolution with the gauge-invariant Hamiltonian should in principle remain within that projection, but truncation and creeping random errors can produce wavefunction components which are redundant gauge copies or worse; a proposal which we find interesting to ameliorate this gauge drift has been put forward by Ball and Cohen [121], and it is to repeatedly project the gauge at short time intervals, and exploit the quantum Zeno effect that blocks the evolution of the measured observable.
@@ 6.1 Time-axial gauge
Here, $\rho_{F}=\psi^{\dagger}T^{a}\psi$ is the fermion color density (taken at each lattice site in the case of the Kogut-Süsskind formulation above). In turn, $\rho_{G}$ is its non Abelian counterpart due to the gluon color charge, and in the lattice formulation, $\rho_{G}=\frac{1}{2}\sum_{\mu}Q_{r\mu}$ is the charge carried by the link that starts at $r$ and ends at $r+\hat{\mu}$.
It is not obvious a priori that any of the states in the discretized “kinematic” Hilbert space actually satisfies Gauss’s condition exactly. Instead, a common procedure is to minimize $\langle\mathcal{G}^{2}\rangle$ over this kinematic space and expect that, in the continuum limit, the minimization leads to the exact satisfaction of Gauss’s law. There are here pitfalls which need to be avoided. For example, if the Fock space/ Hilbert space is defined in terms of normalizable wavefunctions only, as pertains to bound state problems, for example, the limiting procedure could lead to those states satisfying Gauss’s law to fall outside that space. Gauss’s law would then have to be applied, in a weaker sense, to its algebraic dual [124] set (loosely, Dirac bracs) $\langle\psi^{\prime}|$ so that the matrix elements of any observables $\langle\psi^{\prime}|\mathcal{O}\psi\rangle$ would in effect comply with the gauge restriction.
@@ 6.2 Light-front gauge
The light-front gauge is defined by $A^{a+}=0$. With this condition, the classical equations of motion lead to $A^{a-}$ being fixed by a constraint,
$\partial^{+}A^{a-}$ $=$ $2D^{ab\perp}A^{b\perp}-\frac{2}{\partial^{+}}j^{a+}$ (61)
and thus not being quantized, there remaining the two physical degrees of freedom $A^{a\perp}$.
The free term $\partial$ in the covariant derivative $D$ of Eq. (61) is independent of the coupling constant $g$; looking at its purely Yang-Mills part (without the quark color current $j$) it is customary to define $\tilde{A}^{-}$ as
$\tilde{A}^{-}$ $=$ ${1\over\partial^{+}}2\partial^{\perp}A^{\perp}$ (62)
@@
The light-front gauge is well suited for front-form formulations of Hamiltonian dynamics [125] (see [126] for an extended review), where the quantization surface in quantum theory–or the specifications of initial conditions in classical theory–is defined by $x^{+}:=t+z=0$, where $x^{+}$ is referred to as the light-front time. The quantum dynamical evolution then proceeds along $x^{+}$:
@@
Light-front dynamics is useful for several reasons. First, conservation of momentum at interaction vertices, together with the positivity of the front-form momentum $p^{+}=p_{1}^{+}+p_{3}^{+}$ precludes interactions with only creation or only annihilation operators, thereby simplifying the structure of the Hamiltonian [^3] Of course this condition does not apply to only-gluon interactions, since in this case $p^{+}$ can vanish. This exception can, however, be circumvented by employing a regularization procedure that assumes gluons to have an infinitesimally small mass [127, 128].. Relativity requires preserving the commutation relations of the 10 generators of the Poincaré group. In a front form representation 7 of these generators are interaction free, one of them being the boost along $z$ direction, which in this form has a particularly simple structure. The form form is therefore specially convenient for describing processes that involve a preferred direction of motion, as is the case in deep inelastic scattering.
Two main approaches are used to perform numerical calculation in Light-Front QCD, which can be simulated on a quantum computer: Discretized Light-Cone Quantization (DLCQ) [129, 130] and Basis Light-Front Quantization (BLFQ) [131], the later originated from the former. DLCQ discretizes the fundamental fields on the space-time, providing a suitable framework for simulating ab initio fundamental interactions, whereas BLFQ adopts a basis function representation and is well suited for formulating effective theories and computing static hadron observables [132].
@@ 7 Progress towards the Equation of State of Neutron Stars
One of the most anticipated applications of quantum computers is the characterization of many-body systems from the simulation (whether digital or analog) of their basic constituents and interactions. Current research mainly focuses on the calculation of the partition function of canonical ensembles, which depends on temperature $T$, number of particles $N$ and volume $V$ (omitted):
@@
A common method considers the extension of the Hamiltonian parameters $h$ to the complex plane and the characterization of the partition function by its complex zeros (e.g. those of Lee and Yang for the grand-canonical partition function [137, 138] or of Fisher[139] for the canonical one as function of the inverse temperature $\beta$). The behavior of these zeros can also help elucidate the existence of phase transitions: they reside in the complex $\beta$ plane because $Z$ is a sum of positive terms, but as they approach the real axis they can cause nonanalyticities of thermodynamic functions.
@@
$H_{s}=J\left(\sum_{i}\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}\right)+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}$ (65)
@@
The most resource-intensive part of the algorithm is the preparation of the thermal state, whose realization on quantum computers is a field of investigation all by itself. The authors of this extant work [140] use the thermofield double (TFD) method for quantum computers[141]; whose first step is to variationally prepare the state
$\left|\text{TFD}\,(\beta)\right\rangle=\frac{1}{\sqrt{Z(\beta)}}\sum_{n}e^{-\beta E_{n}/2}\left|n\right\rangle_{A}\left|n^{\prime}\right\rangle_{B}$ (68)
in an extended Hilbert space with subsystems $A$ and $B$ with $\left|n\right\rangle$ and $E_{n}$ the eigenstates and eigenenergies of $A$ and $\left|n^{\prime}\right\rangle=\otimes_{j}Y_{j}\left|n\right\rangle$. Subsystem $B$ is then traced out, so that at the end subsystem $A$ is left in the thermal state with density matrix $\rho_{A}=e^{-\beta H_{A}}/Z(\beta)$.
@@
$\partial_{\beta}\left|\Phi(\beta)\right\rangle=\hat{H}\left|\Phi(\beta)\right\rangle$ (69)
@@
converges (at $T=0$) to the ground-state of the Hamiltonian, always provided $\left\langle\Phi(0)\,\right|\left.\Psi_{0}\right\rangle\neq 0$. The implementation on a quantum computer then rests on the ability to efficiently unitarize the non-unitary transformations $e^{-\beta H}$. It has been shown[142] that this can be done provided the Hamiltonian is geometrically k-local, which means that the Hamiltonian can be written as a sum with terms each acting on k-neighbouring qubits $H=\sum_{m}h[m]$. This is expected to happen in Quantum Chromodynamics since the Hamiltonian is a polynomial of degree 4 in the fields, however, some gauges can in principle induce long-range correlations. This needs further case to case study.
A Trotter decomposition can then be used to evolve with each of the $h[m]$ over a small “time” $\Delta\tau$ with $\beta=N\Delta\tau$, $\exp\left\{-\beta\,h[m]\right\}$ is then realized approximately by repeatedly acting with $\exp\left\{-i\Delta\tau A[m]\right\}$, with $A$ a Hermitian matrix decomposed in terms of Pauli strings
$A[m]=\sum_{m}a[m]_{i_{1},...,i_{k}}\sigma_{i_{1}}...\sigma_{i_{k}}=\sum_{m}a[m]_{I}\sigma_{I}$ (71)
acting on the $k$-neigbouring qubits. The coefficients $a[m]$ are then found by solving the linear system $S_{I,I^{\prime}}\,a[m]=b_{I}$ with
$S_{I,I^{\prime}}=\left\langle\Psi\right|\sigma^{\dagger}_{I}\sigma_{I^{\prime}}\left|\Psi\right\rangle,\,\,\,\,b_{I}=\frac{-i}{\sqrt{c}}\left\langle\Psi\right|\sigma^{\dagger}_{I}h[m]\left|\Psi\right\rangle$ (72)
@@
In the context of hadron physics, the Hilbert space dimensionality, the breaking of perturbative expansions and the Monte Carlo sign problems of LQCD are some of the reasons that explain the slow progress towards the characterization of the QCD phase diagram and in particular of the nature of the phase transition between deconfined quark matter and confined hadronic or nuclear media. Such a phase transition is expected to occur at the core of massive Neutron Stars[143]. Thus one can try to derive a suitable EoS by approaching the phase transition from the quark-matter side and to compare it with the astrophysical constraints. Such a system should be described by the Grand-Canonical ensemble, in which the fixed thermodynamic variables are the temperature $T$, the volume $V$ and the chemical potential $\mu$, with partition function:

### v2

@@ 5.1.3 $n$-particle operators
Generalising section 5.1.2, $n$-register operators are constructed by means of projectors (control), $\mathfrak{C}_{ij}=|i\rangle\langle j|$, identities $\mathfrak{i}$ and set/scrap operators $\mathfrak{s}^{\dagger}_{p}/\mathfrak{s}_{p}$. The operator that activates the first register is
@@
$a^{(n)\dagger}_{p,1}{\left|{\Omega}\right\rangle}$ $=$ $\left({\left|{0}\right\rangle}_{P/A}{\left|{0...0}\right\rangle}\right)_{n}...\left({\left|{1}\right\rangle}_{P/A}{\left|{p}\right\rangle}\right)_{1}\ .$ (41)
The most efficient way to implement the symmetrisation for more than two registers is to proceed in a stepwise fashion, assuming previously symmetrised n1n-1 registers and requiring the newly filled register to also be in a symmetric wavefunction with the remaining ones. Hence, the $n$-particle symmetriser [^1] This symmetriser is not idempotent, but satisfies $S^{2}_{n\leftarrow n-1}=\sqrt{n}\ S_{n\leftarrow n-1}$ instead.
@@
$a^{(n)\dagger}_{q}=\sum^{n}_{i=1}a^{(n)\dagger}_{q,i}=\sum^{n}_{i=1}\mathcal{S}_{i\leftarrow(i-1)}\cdot{\text{$\mathbb{P}^{({n-i})}_{{0}}$}}\otimes\left({\text{$\mathfrak{C}_{{1}{0}}$}}\otimes\mathfrak{s}^{\dagger}_{q}\right)_{i}\otimes\,{\text{$\mathbb{P}^{({i-1})}_{{i-1}}$}}\$, (43)
@@
which would be impossible to satisfy with operators acting on single registers (non-trivially) if the memory was not explicitly symmetrised.
These operators satisfy the following commutation relations in the Hilbert subspace of symmetric states
$\left[a^{(n)}_{\rho},a^{(n)\dagger}_{\eta}\right$] $=\delta_{\rho\eta}\left({\text{$\mathfrak{C}_{{0}{0}}$}}\otimes{\text{$\mathfrak{i}$}}\right)_{n}\otimes\mathbb{I}^{(n-1)}\$
$-\ \,S_{n\leftarrow n-1}\cdot\left({\text{$\mathfrak{C}_{{1}{1}}$}}\otimes\mathfrak{s}^{\dagger}_{\rho}\mathfrak{s}_{\eta}\right)_{n}\otimes{\text{$\mathbb{P}^{({n-1})}_{{n-1}}$}}\cdot S_{n\leftarrow n-1}\ .$ (47)
@@
. These definitions can be applied to any operator written in terms of creators and annihilators, for the number operator we have
@@
$=\sum_{j}\,j\,{\text{$\mathbb{P}^{({n})}_{{j}}$}}\$, (48)
where in the last line we assumed a symmetric memory so that $\mathcal{S}_{j\leftarrow j-1}$ (c.f. 42) simplify to $\sqrt{j}$.
@@ 5.2 Fermions
Because the conceptual architecture is very similar, we only briefly touch on anticommuting operators. We change the recursive step-symmetriser operators to step-antisymmetrisers
${\text{$\mathcal{A}_{{n}\leftarrow{n}-1}$}}=\frac{1}{\sqrt{n}}\left({\text{$\mathbb{I}$}}^{\otimes n}-\mathcal{P}_{n(n-1)}-\mathcal{P}_{n(n-2)}-...-\mathcal{P}_{n2}-\mathcal{P}_{n1}\right)\ ,$ (49)
@@
$b^{(n)\dagger}_{q}=\sum^{n}_{j=1}b^{(n)\dagger}_{q,j}=\sum^{n}_{j=1}{\text{$\mathcal{A}_{{j}\leftarrow{j}-1}$}}\cdot{\text{$\mathbb{P}^{({n-j})}_{{0}}$}}\otimes\underbrace{\left({\text{$\mathfrak{C}_{{1}{0}}$}}\otimes\mathfrak{s}^{\dagger}_{q}\right)_{j}}\otimes\,{\text{$\mathbb{P}^{({j-1})}_{{j-1}}$}}\$, (50)
@@
$b^{(n)}_{q}=\sum^{n}_{j=1}b^{(n)}_{q,j}=\sum^{n}_{j=1}{\text{$\mathbb{P}^{({n-j})}_{{0}}$}}\otimes\left({\text{$\mathfrak{C}_{{0}{1}}$}}\otimes{\text{$\mathfrak{s}_{{p}}$}}\right)_{j}\otimes\,{\text{$\mathbb{P}^{({j-1})}_{{j-1}}$}}\cdot\mathcal{A}_{j\leftarrow j-1}\ .$ (51)
They now fulfil the required anticommutation relations
@@ 5.3 Quantum evolution
For interaction terms that change the values of momenta we need to define a discretisation of $p$; for instance, the equidistant grid
$\left\{p_{min},...,p_{-1},p_{0}=0,p_{1},...,p_{max}\right\}\ ,$ (55)
@@
When considering Hamiltonian interaction terms that alter the number of particles, we require a dynamical memory, which changes the number of active registers. Similarly, we can define tadpoles, emission or absorption of particles, etc. For instance, the following operator stands for the evolution associated to the Hamiltonian term corresponding to the quark-gluon vertex
@@ 6 Gauge fixing
The gauge redundancy in Yang-Mills theory or Chromodynamics forces one to select a gauge when fixing the physical states in any Hamiltonian formulation. First and foremost we discuss equal–time quantisation in subsection 6.1. Then we turn to light–front quantisation in subsection 6.2.
With a bird’s eye view, unless one fixes Coulomb gauge which is a complete gauge fixing with an entirely physical Fock space but a convoluted Hamiltonian necessitating approximation [117, 118, 119, 120], the computer will often be approximating a redundant Fock space on which Gauss’s law needs to be employed to project over a physical instance. The latter evolution with the gauge-invariant Hamiltonian should in principle remain within that projection, but truncation and creeping random errors can produce wavefunction components which are redundant gauge copies or worse; a proposal which we find interesting to ameliorate this gauge drift has been put forward by Ball and Cohen [121], and it is to repeatedly project the gauge at short time intervals, and exploit the quantum Zeno effect that blocks the evolution of the measured observable.
@@ 6.1 Time-axial gauge
Here, $\rho_{F}=\psi^{\dagger}T^{a}\psi$ is the fermion colour density (taken at each lattice site in the case of the Kogut-Süsskind formulation above). In turn, $\rho_{G}$ is its non-Abelian counterpart due to the gluon colour charge, and in the lattice formulation, $\rho_{G}=\frac{1}{2}\sum_{\mu}Q_{r\mu}$ is the charge carried by the link that starts at $r$ and ends at $r+\hat{\mu}$.
It is not obvious a priori that any of the states in the discretised “kinematic” Hilbert space actually satisfies Gauss’s condition exactly. Instead, a common procedure is to minimise $\langle\mathcal{G}^{2}\rangle$ over this kinematic space and expect that, in the continuum limit, the minimisation leads to the exact satisfaction of Gauss’s law. There are here pitfalls which need to be avoided. For example, if the Fock space/ Hilbert space is defined in terms of normalizable wavefunctions only, as pertains to bound state problems, for example, the limiting procedure could lead to those states satisfying Gauss’s law to fall outside that space. Gauss’s law would then have to be applied, in a weaker sense, to its algebraic dual [124] set (loosely, Dirac bracs) $\langle\psi^{\prime}|$ so that the matrix elements of any observables $\langle\psi^{\prime}|\mathcal{O}\psi\rangle$ would in effect comply with the gauge restriction.
@@ 6.2 Light-front gauge
The light-front gauge is defined by $A^{a+}=0$. With this condition, the classical equations of motion lead to $A^{a-}$ being fixed by a constraint
$\partial^{+}A^{a-}$ $=$ $2D^{ab\perp}A^{b\perp}-\frac{2}{\partial^{+}}j^{a+}\ ,$ (59)
and thus not being quantised, there remaining the two physical degrees of freedom $A^{a\perp}$.
The free term $\partial$ in the covariant derivative $D$ of Eq. (59) is independent of the coupling constant $g$; looking at its purely Yang-Mills part (without the quark colour current $j$) it is customary to define $\tilde{A}^{-}$ as
$\tilde{A}^{-}$ $=$ ${1\over\partial^{+}}2\partial^{\perp}A^{\perp}\ ,$ (60)
@@
The light-front gauge is well suited for front-form formulations of Hamiltonian dynamics [125] (see [126] for an extended review), where the quantisation surface in quantum theory–or the specifications of initial conditions in classical theory–is defined by $x^{+}:=t+z=0$, where $x^{+}$ is referred to as the light-front time. The quantum dynamical evolution then proceeds along $x^{+}$:
@@
Light-front dynamics is useful for several reasons. First, conservation of momentum at interaction vertices, together with the positivity of the front-form momentum $p^{+}=p_{1}^{+}+p_{3}^{+}$ precludes interactions with only creation or only annihilation operators, thereby simplifying the structure of the Hamiltonian [^2] Of course this condition does not apply to only-gluon interactions, since in this case $p^{+}$ can vanish. This exception can, however, be circumvented by employing a regularisation procedure that assumes gluons to have an infinitesimally small mass [127, 128].. Relativity requires preserving the commutation relations of the 10 generators of the Poincaré group. In a front form representation 7 of these generators are interaction free, one of them being the boost along $z$ direction, which in this form has a particularly simple structure. The front-form is therefore specially convenient for describing processes that involve a preferred direction of motion, as is the case in deep inelastic scattering.
Two main approaches are used to perform numerical calculation in Light-Front QCD, which can be simulated on a quantum computer: Discretised Light-Cone Quantisation (DLCQ) [129, 130] and Basis Light-Front Quantisation (BLFQ) [131], the latter originated from the former. DLCQ discretises the fundamental fields on the space-time, providing a suitable framework for simulating ab initio fundamental interactions, whereas BLFQ adopts a basis function representation and is well suited for formulating effective theories and computing static hadron observables [132].
@@ 7 Progress towards the Equation of State of Neutron Stars
One of the most anticipated applications of quantum computers is the characterisation of many-body systems from the simulation (whether digital or analogue) of their basic constituents and interactions. Current research mainly focuses on the calculation of the partition function of canonical ensembles, which depends on temperature $T$, number of particles $N$ and volume $V$ (omitted):
@@
A common method considers the extension of the Hamiltonian parameters $h$ to the complex plane and the characterisation of the partition function by its complex zeros (e.g. those of Lee and Yang for the grand-canonical partition function [135, 136] or of Fisher[137] for the canonical one as function of the inverse temperature $\beta$). The behavior of these zeros can also help elucidate the existence of phase transitions: they reside in the complex $\beta$ plane because $Z$ is a sum of positive terms, but as they approach the real axis they can cause nonanalyticities of thermodynamic functions.
@@
$H_{s}=J\left(\sum_{i}\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}\right)+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}\ ,$ (63)
@@
The most resource-intensive part of the algorithm is the preparation of the thermal state, whose realisation on quantum computers is a field of investigation all by itself. The authors of this extant work [138] use the thermofield double (TFD) method for quantum computers[139]; whose first step is to variationally prepare the state
$\left|\text{TFD}\,(\beta)\right\rangle=\frac{1}{\sqrt{Z(\beta)}}\sum_{n}e^{-\beta E_{n}/2}\left|n\right\rangle_{A}\left|n^{\prime}\right\rangle_{B}\ ,$ (66)
in an extended Hilbert space with subsystems $A$ and $B$ with $\left|n\right\rangle$ and $E_{n}$ the eigenstates and eigenenergies of $A$ and $\left|n^{\prime}\right\rangle=\otimes_{j}Y_{j}\left|n\right\rangle$. Subsystem $B$ is then traced out, so that at the end subsystem $A$ is left in the thermal state with density matrix $\rho_{A}=e^{-\beta H_{A}}/Z(\beta)$. This procedure has recently been used [140] to prepare thermal states of the Kogut-Sussking Hamiltonian (1+1D QCD, see sec. 4.1) on a chain of trapped-ion qubits. The translational degrees of freedom in the y-direction are entangled with the ion’s internal degrees of freedom storing the qubits in a way that depends on a parameter $\theta_{i}$ for the $i$th ion:
$\left|\psi(\theta_{i})\right\rangle=\cos\left(\theta_{i}/2\right)\left|0,0\right\rangle+\sin\left(\theta_{i}/2\right)\left|1,1\right\rangle\ ,$ (67)
where the computational basis is denoted by $\left|\text{spin},\text{motion}\right\rangle$. Tracing out the translational degrees of freedom, the memory is left in a mixture of bit strings $\left|j\right\rangle\equiv\left|j_{1},j_{2}...j_{N}\right\rangle$ with probabilities $p_{j}(\vec{\theta})=\prod^{N}_{i=1}p_{j_{i}}(\theta_{i})$. This procedure is then used as part of the VQE algorithm to measure the chiral condensate $\hat{\chi}$ over states that minimize the free energy F=ETSF=E-TS. Colour neutrality is then enforced using projection operator $\hat{K}$ to the singlet subspace. The expectation values $\left\langle\hat{O}\right\rangle_{O}$ over singlet subspaces are then defined as
$\left\langle\right.\hat{O}\left.\right\rangle_{O}=\frac{\left\langle\right.\hat{O}\hat{K}\left.\right\rangle}{\left\langle\right.\hat{K}\left.\right\rangle}\$. (68)
@@
$\partial_{\beta}\left|\Phi(\beta)\right\rangle=\hat{H}\left|\Phi(\beta)\right\rangle\ ,$ (69)
@@
converges (at $T=0$) to the ground-state of the Hamiltonian, always provided $\left\langle\Phi(0)\,\right|\left.\Psi_{0}\right\rangle\neq 0$. The implementation on a quantum computer then rests on the ability to efficiently unitarise the non-unitary transformations $e^{-\beta H}$. It has been shown[141] that this can be done provided the Hamiltonian is geometrically k-local, which means that the Hamiltonian can be written as a sum with terms each acting on k-neighbouring qubits $H=\sum_{m}h[m]$. This is expected to happen in Quantum Chromodynamics since the Hamiltonian is a polynomial of degree 4 in the fields, however, some gauges can in principle induce long-range correlations. This needs further case-by-case study.
A Trotter decomposition can then be used to evolve with each of the $h[m]$ over a small “time” $\Delta\tau$ with $\beta=N\Delta\tau$, $\exp\left\{-\beta\,h[m]\right\}$ is then realised approximately by repeatedly acting with $\exp\left\{-i\Delta\tau A[m]\right\}$, with $A$ a Hermitian matrix decomposed in terms of Pauli strings
$A[m]=\sum_{m}a[m]_{i_{1},...,i_{k}}\sigma_{i_{1}}...\sigma_{i_{k}}=\sum_{m}a[m]_{I}\sigma_{I}\ ,$ (71)
acting on the $k$-neighbouring qubits. The coefficients $a[m]$ are then found by solving the linear system $S_{I,I^{\prime}}\,a[m]=b_{I}$ with
$S_{I,I^{\prime}}=\left\langle\Psi\right|\sigma^{\dagger}_{I}\sigma_{I^{\prime}}\left|\Psi\right\rangle\ ,\,\,\,\,b_{I}=\frac{-i}{\sqrt{c}}\left\langle\Psi\right|\sigma^{\dagger}_{I}h[m]\left|\Psi\right\rangle\ ,$ (72)
@@
In the context of hadron physics, the Hilbert space dimensionality, the breaking of perturbative expansions and the Monte Carlo sign problems of LQCD are some of the reasons that explain the slow progress towards the characterisation of the QCD phase diagram and in particular of the nature of the phase transition between deconfined quark matter and confined hadronic or nuclear media. Such a phase transition is expected to occur at the core of massive Neutron Stars[142]. Thus one can try to derive a suitable EoS by approaching the phase transition from the quark-matter side and to compare it with the astrophysical constraints. Such a system should be described by the Grand-Canonical ensemble, in which the fixed thermodynamic variables are the temperature $T$, the volume $V$ and the chemical potential $\mu$, with partition function: